# A Survey of Large Language Models in Psychotherapy: Current Landscape and Future Directions

### v1

@@ 3.2 Diagnosis
is based on a fixed set of data, typically derived from complete dialogues or social media posts. Galatzer-Levy et al. (2023) highlighted the effectiveness of Med-PaLM 2 in psychiatric condition assessment from patient interviews and clinical descriptions without specialized training. Similarly, Jiang et al. (2024b) showcased LLMs’ superior performance on depression and anxiety detection on Russian datasets, particularly with noisy or small datasets. Hengle et al. (2024) evaluated PLMs and LLMs on multi-label classification in depression and anxiety, underscoring the ongoing challenges in applying LMs to mental health diagnostics. Besides, Lan et al. (2024c) introduced DORIS, a depression detection system integrating text embeddings with LLMs, utilizing symptom features, post-history, and mood course representations to make diagnostic predictions and generate explanatory outputs. Kuzmin et al. (2024) developed ADOS-Copilot for ASD diagnosis through diagnostic dialogues, employing In-context Enhancement, Interpretability Augmentation, and Adaptive Fusion based on real-world ADOS-2 clinic scenarios.
@@ 3.3 Treatment
refrains from providing a holistic therapy role but instead offers targeted support such as rewriting suboptimal counselor responses, generating controlled reappraisal prompts, or aiding clinicians in specific tasks. For example, Welivita and Pu (2023) proposed to rewrite responses that violate MI principles into MI-adherent forms, ensuring more consistent therapeutic dialogue. Meanwhile, Sharma et al. (2023) and Maddela et al. (2023) focused on generating single-turn reframes of negative thoughts—often anchored in cognitive distortions—through controlled language attributes. On the detection side, Moon and Bhattacharyya (2024) built a multimodal pipeline to identify depression and provide CBT-style replies, albeit with an emphasis on technological assistance rather than full-fledged therapy. In the Chinese context, Lin et al. (2024) combined cognitive distortion detection with “positive reconstruction,” demonstrating a single-round rewrite approach for negative or distorted statements, while Na (2024) showcased a structured Q&A format that offers professional yet succinct CBT-based responses. From a knowledge-distillation angle, Brown et al. (2024) demonstrated how smaller models could replicate GPT-4’s MI-style reflective statements, and Zhan et al. (2024) introduced a lighter-weight framework RESORT to guide smaller LLMs toward effective cognitive reappraisal prompts, thus enabling broader accessibility of self-help tools.
@@
pivots toward generating synthetic yet realistic patient behaviors or multi-level feedback to train or support mental health practitioners. For instance, Chaszczewicz et al. (2024) leveraged LLMs to deliver multi-tier feedback on novice peer counselors’ conversational skills, significantly reducing the need for continuous expert oversight. Similarly, Wang et al. (2024b) using LLM-driven patient simulations that help trainees practice CBT core skills in a controlled, repeatable setup. In the realm of assessing therapy quality, Yosef et al. (2024) showcased a digital patient system to evaluate MI sessions, employing AI-generated transcripts to differentiate novice, intermediate, and expert therapeutic skill levels. Complementarily, Louie et al. (2024) offered Roleplay-doh, a pipeline wherein domain experts craft specialized principles that guide LLM-based role-playing agents, thereby providing customizable training for new therapists.
@@
targets the appraisal of therapy dialogue, counselor techniques, and treatment processes, typically without delivering direct interventions to clients. For instance, Lee et al. (2024e) augmented crisis counseling outcome prediction by fusing annotated counseling strategies with LLM-derived features, achieving substantially improved accuracy. In the Chinese context, Zhang et al. (2024a) introduced CPsyCoun, employing reports-based dialogue reconstruction and automated evaluation to verify counseling realism and professionalism. Beyond single-session analyses, Wang et al. (2024a) used simulated clients to assess perceived therapy outcomes, while Chiu et al. (2024) created the BOLT framework for systematically comparing LLM-based therapy behaviors with high- and low-quality human sessions. Further extending to online counseling, Li et al. (2024a) proposed an LLM-based approach to measure therapeutic alliance, whereas Shapira and Alfi-Yogev (2024) delineated therapist self-disclosure classification as a new NLP task. In the MI domain, Sun et al. (2024) and Cohen et al. (2024) collected bilingual transcripts to systematically annotate therapist–client exchanges for behavior coding and global scores, respectively. Additionally, multi-session perspectives emerge in Na et al. (2024), who proposed IPAEval to track long-term progress from the client’s viewpoint, and Nguyen et al. (2024) analyzed conversation redirection and its impact on patient–therapist alliance over multiple sessions. Finally, Iftikhar et al. (2024) and Zhang et al. (2024b) explored the disparities between LLM- and human-led CBT sessions, highlighting gaps such as empathy and cultural nuance while also introducing CBT-Bench to probe LLMs’ deeper psychotherapeutic competencies.
@@ 4 Current Landscape
Figure 3: Distribution analysis of the current landscape.
@@ 5 Future Directions
Integrative Psychotherapy Framework.
While many existing studies focus on a single dimension of psychotherapy, real-world practice involves a continuous process that spans assessment, diagnosis, and intervention Waszczuk et al. (2017). Moreover, these stages typically unfold over multiple sessions, necessitating iterative, multi-turn interactions that incorporate the evolving context of each patient. Future work could therefore aim to develop an end-to-end conversational framework that seamlessly spans from initial evaluation to personalized intervention. By maintaining a system grounded on ongoing, context-sensitive engagement, models could dynamically update assessments and diagnoses over time, ultimately providing more responsive and individualized care.
Addressing Evolving and Multifaceted Nature of Psychotherapy.
Psychotherapy commonly involves shifting symptoms, comorbidities, and nuanced patient experiences, making static or single-label predictions insufficient. Models should integrate multi-label and temporal data to capture how symptoms and emotional states evolve, while avoiding the pitfalls of incomplete symptom detection. For instance, focusing solely on the depressive features of a bipolar patient could lead to an inaccurate diagnosis if the manic phase is overlooked Lee et al. (2024a). Furthermore, current research suggests that LLMs often struggle with multi-label tasks hee So et al. (2024); Mohammadi et al. (2024), highlighting the need for improved model architectures and algorithms that can better account for these complexities.
Resource Infrastructure and Open-Source Tools.
Current research heavily relies on commercial closed-source models, lacking reproducible open-source evaluation methods and multilingual data. Notably, developing multilingual datasets should not solely rely on translating English resources, as psychological research indicates that cultural context plays a critical role in mental health. English-based translations cannot fully substitute for culturally specific data from other languages Watters (2010); Abdelkadir et al. (2024).
@@
Most studies to date have concentrated on prevalent conditions such as depression and anxiety, leaving complex or less common disorders underexplored. Additionally, research tends to focus on a limited range of therapeutic modalities—primarily cognitive behavioral therapy. Future work could broaden both the range of disorders and the variety of therapeutic approaches, such as humanistic Schneider and Krug (2010) and dialectical behavior therapy Lynch et al. (2006), to better reflect clinical realities Norcross et al. (2022). Such an expansion could deepen the theoretical underpinnings of LLM-based psychotherapy tools and enhance the quality and relevance of digital interventions.
@@ 6 Conclusion
LLMs hold significant promise for revolutionizing psychotherapy by enhancing assessment, diagnosis, and treatment processes through dynamic, context-sensitive interactions. Despite the progress made, key challenges such as linguistic biases, limited disorder coverage, and underrepresented therapeutic models persist. Future research should focus on creating integrative, multi-turn systems that span the entire psychotherapy process while addressing the evolving nature of mental health conditions. Expanding resources, embracing diverse therapeutic approaches, and improving model architectures will be crucial in making LLM-driven psychotherapy tools more effective, inclusive, and adaptable.
@@
This survey paper, while comprehensive for LLMs in psychotherapy, has several limitations: 1) The studies reviewed primarily focus on the application of LLMs in psychotherapy, and there may be relevant research in adjacent fields or interdisciplinary domains that was not included. 2) Due to the rapidly evolving nature of this area, some recent advancements may not be captured. The scope of this survey is limited to the available literature and may overlook emerging trends or unpublished findings. 3) The review primarily examines studies in English, which could introduce a bias towards research from English-speaking countries, potentially overlooking important cultural perspectives. 4) While we provide a taxonomy of LLM applications in psychotherapy, this framework may not fully encompass the complexity of real-world clinical settings or the diverse range of therapeutic approaches currently in practice.

### v2

@@ 3.2 Diagnosis
is based on a fixed set of data, typically derived from complete dialogues or social media posts. Galatzer-Levy et al. (2023) highlighted the effectiveness of Med-PaLM 2 in psychiatric condition assessment from patient interviews and clinical descriptions without specialized training. Similarly, Jiang et al. (2024c) showcased LLMs’ superior performance on depression and anxiety detection on Russian datasets, particularly with noisy or small datasets. Hengle et al. (2024) evaluated PLMs and LLMs on multi-label classification in depression and anxiety, underscoring the ongoing challenges in applying LMs to mental health diagnostics. Besides, Lan et al. (2024c) introduced DORIS, a depression detection system integrating text embeddings with LLMs, utilizing symptom features, post-history, and mood course representations to make diagnostic predictions and generate explanatory outputs. Kuzmin et al. (2024) developed ADOS-Copilot for ASD diagnosis through diagnostic dialogues, employing In-context Enhancement, Interpretability Augmentation, and Adaptive Fusion based on real-world ADOS-2 clinic scenarios.
@@ 3.3 Treatment
refrains from providing a holistic therapy role but instead offers targeted support such as rewriting suboptimal counselor responses, generating controlled reappraisal prompts, or aiding clinicians in specific tasks. For example, Welivita and Pu (2023) proposed to rewrite responses that violate MI principles into MI-adherent forms, ensuring more consistent therapeutic dialogue. Meanwhile, Sharma et al. (2023) and Maddela et al. (2023) focused on generating single-turn reframes of negative thoughts–often anchored in cognitive distortions–through controlled language attributes. On the detection side, Moon and Bhattacharyya (2024) built a multimodal pipeline to identify depression and provide CBT-style replies, albeit with an emphasis on technological assistance rather than full-fledged therapy. In the Chinese context, 43 combined cognitive distortion detection with “positive reconstruction,” demonstrating a single-round rewrite approach for negative or distorted statements, while Na (2024) showcased a structured Q&A format that offers professional yet succinct CBT-based responses. From a knowledge-distillation angle, Brown et al. (2024) demonstrated how smaller models could replicate GPT-4’s MI-style reflective statements, and Zhan et al. introduced a lighter-weight framework RESORT to guide smaller LLMs toward effective cognitive reappraisal prompts, thus enabling broader accessibility of self-help tools.
@@
pivots toward generating synthetic yet realistic patient behaviors or multi-level feedback to train or support mental health practitioners. For instance, Chaszczewicz et al. (2024) leveraged LLMs to deliver multi-tier feedback on novice peer counselors’ conversational skills, significantly reducing the need for continuous expert oversight. Similarly, Wang et al. (2024b) introduced LLM-driven patient simulations that help trainees practice CBT core skills in a controlled, repeatable setup. In the realm of assessing therapy quality, Yosef et al. (2024) showcased a digital patient system to evaluate MI sessions, employing AI-generated transcripts to differentiate novice, intermediate, and expert therapeutic skill levels. Complementarily, Louie et al. (2024) offered Roleplay-doh, a pipeline wherein domain experts craft specialized principles that guide LLM-based role-playing agents, thereby providing customizable training for new therapists.
Figure 3: Distribution Analysis of The Current Landscape. Panel (a) indicates English is the predominant language (70%), with Chinese also represented (24.2%). Panel (b) shows that 32% of studies address mental disorders, with Depression and Anxiety being the most common topics within this group. Panel (c) reveals that 32.8% of studies incorporate psychotherapy theories, where CBT is the most frequently applied.
@@
targets the appraisal of therapy dialogue, counselor techniques, and treatment processes, typically without delivering direct interventions to clients. For instance, Lee et al. (2024e) augmented crisis counseling outcome prediction by fusing annotated counseling strategies with LLM-derived features, achieving substantially improved accuracy. In the Chinese context, Zhang et al. (2024) introduced CPsyCoun, employing reports-based dialogue reconstruction and automated evaluation to verify counseling realism and professionalism. Beyond single-session analyses, Wang et al. (2024a) used simulated clients to assess perceived therapy outcomes, while Chiu et al. (2024) created the BOLT framework for systematically comparing LLM-based therapy behaviors with high- and low-quality human sessions. Further extending to online counseling, Li et al. (2024a) proposed an LLM-based approach to measure therapeutic alliance, whereas Shapira and Alfi-Yogev (2024) delineated therapist self-disclosure classification as a new NLP task. In the MI domain, Sun et al. (2024) and Cohen et al. (2024) collected bilingual transcripts to systematically annotate therapist–client exchanges for behavior coding and global scores, respectively. Additionally, multi-session perspectives emerge in Na et al. (2024), who proposed IPAEval to track long-term progress from the client’s viewpoint, and Nguyen et al. (2024) analyzed conversation redirection and its impact on patient–therapist alliance over multiple sessions. Finally, Iftikhar et al. (2024) and Zhang et al. (2025) explored the disparities between LLM- and human-led CBT sessions, highlighting gaps such as empathy and cultural nuance while also introducing CBT-Bench to probe LLMs’ deeper psychotherapeutic competencies.
@@ 4 Current Landscape
4.1 Overview
@@
4.2 Fragmented Approaches
Traditionally, LLM-based psychotherapy tools have addressed assessment, diagnosis, and treatment separately. Recently, a few studies have started to explore integrative approaches spanning multiple stages. Despite these emerging integrative efforts, the systems remain limited, typically addressing only two stages without achieving full continuity. Additionally, fragmentation occurs not only across these three dimensions but also at more granular levels; for example, some methodologies are narrowly focused on assessing single disorders Tu et al. (2024); Bao et al. (2024), further limiting their applicability and integration potential in broader psychotherapy contexts.
4.3 Critical Issues and Risks
Dynamic Symptom Representation.
Psychotherapy commonly involves shifting symptoms, comorbidities, and nuanced patient experiences, making static or single-label predictions insufficient. Current models fail to adequately capture multi-label conditions and temporal symptom fluctuations, leading to incomplete or inaccurate assessments Lee et al. (2024a).
Linguistic Resource Bias and Translation Critique.
Most psychotherapy-oriented LLMs are trained primarily on English datasets, with some researchers attempting to expand linguistic coverage through translation. However, recent studies highlight significant cultural specificity in mental health disorders Watters (2010); Abdelkadir et al. (2024), making direct translation of datasets unreliable for accurately capturing psychological nuances across different cultures.
Diagnostic Risks.
Current approaches to automated diagnosis often struggle to gain acceptance among clinical practitioners due to concerns about reliability and patient safety. Despite the psychology community increasingly favoring transdiagnostic methods Dalgleish et al. (2020), a segment of NLP researchers continues to emphasize diagnosis-specific studies, indicating a notable divergence in research priorities.
@@ 5 Future Directions
Continuous Multi-Stage Modeling.
Psychotherapy inherently involves continuous interactions that progress through assessment, diagnosis, and intervention phases. Existing research indicates that several leading foundation models exhibit negligible hallucination issues in the medical domain Kim et al. (2025), providing a promising foundation for integrating these stages, as minimizing hallucinations is crucial for ensuring the accuracy and reliability required for continuous patient state tracking across psychotherapy stages. Future models should aim for an evolving representation of patient states, ensuring consistency and coherence across the entire therapeutic process rather than handling stages as isolated segments.
Real-Time Adaptability Grounded in Psychological Theory.
The development of real-time adaptive strategies represents a significant step beyond current static models. Current technical advancements, such as retrieval-augmented generation and long-context memory techniques Jo et al. (2024), provide the necessary technical foundations for such strategies. Instead of simply reproducing patterns found in pre-collected dialogue datasets, future LLM applications should incorporate these advanced contextual memory mechanisms informed by established psychological theories. Such systems would dynamically interpret patient cues, adjusting interventions immediately in response to subtle shifts in emotional and cognitive states. This approach significantly surpasses mere language-style mimicry achieved through simple fine-tuning on existing datasets, enabling deeper, theoretically-informed therapeutic engagement.
@@
Future research should prioritize diversification in terms of both disorders addressed and therapeutic methodologies employed. The current focus on common disorders like anxiety and depression has led to an imbalanced research landscape. There is a pressing need to incorporate underrepresented conditions, such as bipolar disorder and personality disorders, alongside a broader spectrum of psychotherapeutic frameworks, including psychodynamic Shedler (2010) and existential-humanistic approaches Schneider and Krug (2010). Such expansion would help address existing blind spots, contributing to a more inclusive and comprehensive application of LLMs in psychotherapy.
@@ 6 Conclusion
LLMs hold significant promise for revolutionizing psychotherapy by enhancing assessment, diagnosis, and treatment. However, the current landscape reveals critical limitations: research is often fragmented across these stages, exhibits notable biases in linguistic and disorder coverage, and underutilizes diverse psychotherapeutic theories. To overcome these challenges, future work must focus on creating continuous, multi-stage models that are grounded in psychological theory and capable of real-time adaptation to evolving patient states. Expanding the scope of addressed disorders and therapeutic modalities will be crucial for developing LLM-driven psychotherapy tools.
@@
We remind the readers that this survey has the following limitations: 1) The studies reviewed primarily focus on the application of LLMs in psychotherapy, and there may be relevant research in adjacent fields or interdisciplinary domains that was not included. 2) Due to the rapidly evolving nature of this area, some recent advancements may not be captured. The scope of this survey is limited to the available literature and may overlook emerging trends or unpublished findings. 3) While we provide a taxonomy of LLM applications in psychotherapy, this framework may not fully encompass the complexity of real-world clinical settings or the diverse range of therapeutic approaches currently in practice.